OpenStudy (anonymous):

Please, can anyone tell me where and how to find the answer regarding my today's question?

Directrix (directrix):

What is today's question?

OpenStudy (anonymous):

2 questions 1) solve the system of equations dx/dt=4x+3y and dy/dt=-x+t with x(0)=2, y(0)=-1 2) Find the surface area of that portion of the surface z=1- square-root of x^2+y^2 that lies in the first octant. thanks

OpenStudy (loser66):

is it differential equation? @primeralph what do you think?

OpenStudy (primeralph):

@Loser66 Yeah.

OpenStudy (primeralph):